OpenStudy (anonymous):

Which of the following is true about the graph of f(x)=x-2/x^2-2x Vertical Asymptote at x = 0; Slant Asymptote at y = x Vertical Asymptote at x = 0; Horizontal Asymptote at y = 1 Vertical Asymptote at x = 0; Hole at x = 2; Horizontal Asymptote at y = 1 Vertical Asymptote at x = 0; Hole at x = 2; Horizontal Asymptote at y = 0

OpenStudy (campbell_st):

well if you factorise the denominator you get \[f(x) = \frac{x -2}{x(x - 2)}\] which can be simplifed to f(x) = 1/x so the curve has a vertical asymptote at x = 0 and horizontal asymptote at y = 0 by x cannot equal 2 since it will result in a zero denominator... in the original function but since (x -2) can be cancels then x = 2 is a point of discontinuity. (a Hole)

OpenStudy (anonymous):

so...whats the answer?

OpenStudy (campbell_st):

did you read the information..? as the answer is in there...